OpenStudy (anonymous):

In the right triangle, find the length of the side not given. Give an exact answer AND an approximation to three decimal places. a=8 , b=10

OpenStudy (anonymous):

is c the hypotnenuse? if so then use \[c^2=a^2+b^2\] \[c^2=8^2+10^2\] \[c^2=64+100=164\] \[c=\sqrt{164}\]

OpenStudy (anonymous):

simplest radical form is \[2\sqrt{41}\] and decimal is 12.806 from a calculator